A ball is thrown vertically upward and returns to its starting point 20 seconds later. Find its a) initial velocity; b) the maximum height it has reached.

 



1
Expert's answer
2021-09-26T18:47:46-0400

Given:

"t=20\:\rm s"


The time of flight

"t=\frac{2v_0}{g}"

Hence,

a) initial velocity

"v_0=gt/2=9.8*20/2=98\:\rm m/s"

b) the maximum height that ball reached

"h_{\max}=\frac{v_0^2}{2g}=\frac{98^2}{2*9.8}=490\:\rm m"